OpenStudy (anonymous):

Evaluate the function for the given value of x. f(x) = -6x2 - 4x - 4 , x = - 2 A. f( - 2) = -36 B. f( - 2) = -17 C. f( - 2) = -20 D. f( - 2) = -28

OpenStudy (anonymous):

just plug in x=-2

OpenStudy (anonymous):

f(x) = -6x2 - 4x - 4 f(-2) = -6(-2)^2 - 4(-2) - 4

OpenStudy (anonymous):

Now simplify

OpenStudy (anonymous):

\[f(-2) = -6(-2)^2 - 4(-2) - 4\]

OpenStudy (anonymous):

If you have f(x), then saying x = -2 means the same thing as f(-2). \[f(x) = -6x^{2} - 4x - 4\]\[f(-2) = -6(-2)^{2} - 4(-2) - 4\]

OpenStudy (anonymous):

I'll simplify it a little bit more. Try to solve it after this. \[f(-2) = -6(-2)^{2} - 4(-2) - 4\]\[f(-2) = -6(4) + 8 - 4\]\[f(-2) = ?\]

OpenStudy (anonymous):

Do you know how to multiply and add?

OpenStudy (anonymous):

\[f(-2)=-6*4-8-4\] [f(-2)=-6*(-2)^2-4*(-2)-4\] \[f(-2)=-24-12\]\[f(-2)=-36\]

OpenStudy (anonymous):

@muhammad9t5 +8 -4= 4 and not -12 :)

OpenStudy (anonymous):

Oh sorry I didn't even see this, I got a -36 is this right. Sorry I was getting a lil depressed over something.

OpenStudy (anonymous):

@samyflashy its negative 8 not the positive 8.

OpenStudy (anonymous):

so then D ? I'm sorry bare with me lol

OpenStudy (anonymous):

I'm pretty convinced its A.

OpenStudy (anonymous):

C is right answer.